f (x) be defined as follows


 f(x) = {x^ 2+8x+15/x+3 ifx<−3

x^2 − 7 if x ≥ −3 }


Prove from first principles (i.e.an ε − δ proof) that f is continuous at the point x = −3.


1
Expert's answer
2021-05-12T07:00:18-0400

"f(x) = \dfrac{x^2+8x+15}{x+3}" if, "x<-3"


"= x^2-7" if, "x\ge -3"


"RHL = limx \rightarrow-3^{+}(x^2-7) = (-3)^2-7 = 2"


"LHL = limx \rightarrow-3^{-} \dfrac{x^2+8x+15}{x+3}"


"= limx \rightarrow -3^{-}\dfrac{(x+3)(x+5)}{(x+3)}"


"= limx \rightarrow -3^{-}(x+5)"


"= -3+5 = 2"

Also f(x) at "x = -3 = (-3)^2-7 = 9-7 = 2"


We can clearly see that,


"LHL=RHL=f(-3)" Hence the given function is continuous at "x=-3."